HUBBLE SPACE TELESCOPE DAILY REPORT #2221 PERIOD COVERED: 0000Z (UTC) 10/01/98 - 0000Z (UTC) 10/02/98 Daily Status Report as of 275/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of WF/PC-2 7712 (WF/PC-2 Cycle 7 Supplemental Darks pt2/3) The WF/PC-2 was used to make a series of Cycle 7 supplemental calibration darks. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed WF/PC-2/STIS/MA2/NIC/1/2/3 8073 (HDF-South Observations) The WF/PC-2, The Space Telescope Imaging Spectrograph (MA2) and the Near Infrared Camera (Cameras One, Two and Three) all were used to make deep observations of the Hubble Deep fields- South (HDFS-22333). This program is designed to provide deep observations of a high-galactic latitude field in the southern continuous viewing zone. The STIS field is centered on a bright QSO at z=2.25. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 4 Successful : 4 Scheduled Reacquisitions : 9 Successful : 9 No Loss of Lock. 2.2 FHST Updates: Scheduled : 4 Successful : 4 2.3 Operations Notes: The Hubble Deep Field-South (HDF-S) campaign continues as scheduled. There have been no significant problems reported. Using ROP SR-1, the SSR EDAC error counter was reset three times. The FGS-2 Temperature EFG-6T went out of limits for several observations from 275/0553Z through 0630Z. HSTAR 6700 was written. At 275/0835Z NPYTILTP (NICMOS PAM Tilt Position) was flagged out-of-limit for one sample. There were no NICMOS transitions at that time HST was coming out of the SAA. HSTAR 6701 was written. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH